Your firm currently has 1,543,589 shares of stock outstanding that sell for 18.35. Additionally, you've issued 23,052 bonds, with a par value of S1,000. They currently sell for 94.1% of par. What is the firm's equity weight? (Enter your answer as a percentage to 1 decimal place. ex: 12.3)
Answer #1

Total value of:

Equity=1,543,589*18.35=$28324858.2

Bonds=23,052*(1000*94.1%)=$21691932

Total value=(28324858.2+21691932)=$50016790.2

Hence equity weight=(28324858.2/50016790.2)

=56.6%(Approx).